# HyperBlocker: Accelerating Rule-based Blocking in Entity Resolution using GPUs

### v1

@@ 7. Related Work
However, EPG differs from existing query optimizations: (1) EPG optimizes the execution, no matter what comparisons (e.g., equality or similarity) are adopted, while many DBMS optimizers struggle when similarity comparisons are encoded as UDFs, e.g., SQL Server (sql, 2024) restricts UDFs to a single thread, and PostgreSQL (pos, 2024) treats UDFs as black boxes. This said, EPG solves a more specialized problem, beyond general query optimization, for supporting arbitrary comparisons. (2) It is hard for most optimizers to accurately estimate the runtime performance of UDFs (Rheinländer et al., 2017), which may depend on specific measures, thresholds, and data (i.e., data-awareness), while we consider both the time and the selectivity of predicates, using a learned and LSH-based model for accurate estimation. (3) EPG employs tree structures and bitmaps, to effectively handle the disjunction logic behind blocking and to reuse computation, while traditional DBMS may be forced to perform full scan when evaluating OR operations. (4) EPG also produces a data partitioning scheme based on the execution tree as a by-product, to coordinate across multiple GPUs. More in-depth analysis about EPG with the plans generated with DMBS and UDFs optimizations can be found in (ful, 2024).
@@ 8. Conclusion
There are some future topics: (a) accelerate the evaluation of multi-variable rules, e.g.,  $\mathsf{REEs}$ (Fan et al., 2022), using GPUs; (b) give a different plan on each partition; (c) explore the materialization of partial evaluation results to avoid divergence and (d) investigate whether EPG and traditional optimizers can complement/enhance each other.

### v2

@@ 7. Related Work
EPG differs from existing query optimizations: (1) EPG optimizes the execution, no matter what comparisons (e.g., equality or similarity) are adopted, while many DBMS optimizers struggle when similarity comparisons are encoded as UDFs, e.g., SQL Server (sql, 2024) restricts UDFs to a single thread, and PostgreSQL (pos, 2024) treats UDFs as black boxes. This said, EPG solves a more specialized problem, beyond general query optimization, for arbitrary comparisons. (2) It is hard for most optimizers to accurately estimate the runtime performance of UDFs (Rheinländer et al., 2017), which may depend on specific measures/data, while we consider the time/selectivity of predicates, using learned and LSH-based models for accurate estimation. (3) EPG employs tree structures and bitmaps, to effectively handle the disjunction logic behind blocking and to reuse computation, while traditional DBMS may be forced to perform full scan when evaluating OR operations. (4) EPG produces a data partitioning scheme based on the execution tree as a by-product, to coordinate across multiple GPUs.
@@ 8. Conclusion
There are some future topics: (a) give a different plan on each partition; (b) explore the materialization of partial evaluation results to avoid divergence and (c) investigate whether EPG and traditional optimizers can complement/enhance each other.
Acknowledgements.
We sincerely thank Wenfei Fan, Shuhao Liu, Yaoshu Wang, and Weijie Ou for their comments and helpful discussions.